Spiders may "tune" strands of their webs to give enhanced response at frequencies corresponding to the frequencies at which desirable prey might struggle. Orb web silk has a typical diameter of 0.0020 mm, and spider silk has a density of 1300 kg/m^3.

To give a resonance at 130 Hz , to what tension must a spider adjust a 17 cm long strand of silk?

Express your answer to two significant figures and include the appropriate units.
Answer #1

f = v/2L

v = sqrt (T / u)

where u = m/L

combining above two equations

2fL = sqrt (T / u)

squaring both sides

4f2L2 = T / u

4f2L2 = T*L / m

4f2L = T/m

T = 4f2Lm

so,

m = density * volume

m = 1300 * 0.17 * pi * (1e-6)2 = 6.9429e-10 Kg

so,

T = 4*1302 * 0.17 * 6.9429e-10

T = 7.978e-6 N